OpenStudy (anonymous):

The perimeter of a triangle is 7a - 11b. If two of its sides are 2a + b and a - 9b, what is the third side?

OpenStudy (anonymous):

perimeter means all the lenghts of the sides added together. so you know the two sides and the perimeter so you just have to subtract the length of the two sides from the perimeter

OpenStudy (saifoo.khan):

The perimeter of a triangle is the total length of all three sides. P = x + y + z We're told that the perimeter is 7a - 11b, one leg (x) is 2a + b, and another leg (y) is a - 9b. Plug these in, then solve for z, the length of the third side: 7a - 11b = (2a + b) + (a - 9b) + z 7a - 11b = 3a - 8b + z 4a - 3b = z The third leg is 4a - 3b.

OpenStudy (saifoo.khan):

thnk-ku! :)

OpenStudy (anonymous):

What must be added to 3x -7 to make x^2 + 4x -1?

OpenStudy (saifoo.khan):

im not sure about it.. x^2 + x +6

OpenStudy (saifoo.khan):

what grade math is that?

OpenStudy (anonymous):

Perform the operation indicated: (5m^3 + 3mn^2) + (-7m^2n + mn^2 - n^3) - (mn^2 + 2m^3)

OpenStudy (anonymous):

alg 2

OpenStudy (saifoo.khan):

\[3m^3 + 3mn^2 -7m^2n -n^3 \]

OpenStudy (saifoo.khan):

lol, that was Lonnggg.

OpenStudy (anonymous):

lol your my hero:)

OpenStudy (saifoo.khan):

ty! :)

OpenStudy (anonymous):

check it it says it wrong:(

OpenStudy (saifoo.khan):

5m^3 + 3mn^2 -7m^2n -n^3 -2m^3

OpenStudy (saifoo.khan):

u inserted this right? :O

OpenStudy (anonymous):

ya i did :(